Solve 4 / 3 – 1 / 2 Solve 4 / 3 – 1 / 2 Solve 4 / 3 – 1 / 2

Solve
4 / 3 – 1 / 2
Solve
4 / 3 – 1 / 2
Solve
4 / 3 – 1 / 2

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

4 / 3 – 1 / 2
Taking LCM 6
= [(4 × 2)–(1 × 3)] / 6
= (8–3) /6
= 5 / 6

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
4/3 - 1/2
 
Taking LCM as 6,
(4×2)-(1×3)/ 6
= (8 - 3)/6
= 5/6
 
Hope it helps!
Thank you ✌️